A car journey of 300km lasts 4 hours. Parts of this journey is on a freeway at an average speed of 90 km/h. The rest is on country roads at an average speed of 70 km/h. Let T be the time (in hours) spent on the freeway.
I've worked out that in terms of T, the number of hours travelling on country roads is 4-T. And the distance travelled on the freeway is 90T and on the country roads is 70T.
The final part of the question is: Find T and the distance travelled on each type of road.
Thanks in advance to anyone who can help me.

Answer by mananth(16946)

A car journey of 300km lasts 4 hours. Parts of this journey is on a freeway at an average speed of 90 km/h. The rest is on country roads at an average speed of 70 km/h. Let T be the time (in hours) spent on the freeway.
I've worked out that in terms of T, the number of hours travelling on country roads is 4-T. And the distance travelled on the freeway is 90T and on the country roads is 70T.
The final part of the question is: Find T and the distance travelled on each type of road.
..
let distance on freeway be x km
distance on country road = 300-x
...
freeway speed = 90 kmph
country speed = 70kmph.
..
t=d/r
...
time country + time freeway = total time = 4 hours
x/90 + (300-x)/70 = 4
...
(70x+90(300-x))/6300 =4
(70x+27,000-90x=4*6300
-20x=-27000+25200
-20x=-1800
/-20
x= 90 km. freeway distance
freeway time = d/r= 90/90 = 1 hour
...
country distance = 300-90 = 210 km
country time = 210/70 = 3 hours
...
m.ananth@hotmail.ca